# Some lemmas on spectral radius of graphs: including an application

### v2

@@ 2 Preliminaries
Now consider $g(x)=8.5+\frac{7.5}{x-1}+\frac{16}{x-4}-x$ for $x\geq\sqrt{130}$. Clearly, $g(x)$ is decreasing for $x\geq\sqrt{130}$. By a calculation, $g(\sqrt{130})=-0.01\cdots<0$. Thus $g(x)<0$ for any $x\geq\sqrt{130}$. Then $\sum^{\infty}_{i=1}(\frac{W^{i}(G^{1})}{\rho_{1}^{i}}-\frac{W^{i}(G^{2})}{\rho_{1}^{i}})<0$, or equivalently, $\sum^{\infty}_{i=1}\frac{W^{i}(G^{1})}{\rho_{1}^{i}}<\sum^{\infty}_{i=1}\frac{W^{i}(G^{2})}{\rho_{1}^{i}}$. It follows that $f(G_{1},\rho_{1})>f(G_{2},\rho_{1})$. Note that $f(G_{1},\rho_{1})=1$. So, $f(G_{2},\rho_{1})<1$. Recall that $f(G_{2},x)$ is strictly increasing in its domain, and $f(G_{2},\rho_{2})=1$. We must have $\rho_{1}<\rho_{2}$. This completes the proof. $\Box$
@@
$\rho_{j}\geq\rho(K_{1}\vee\overline{K_{|T|+\sum_{1\leq i\leq 5}n_{i}}})=\sqrt{|T|+\sum_{1\leq i\leq 5}n_{i}}\geq\max\left\{\sqrt{310},\sqrt{\sum_{1\leq i\leq 5}n_{i}}\right\}.$
@@
$\sum^{\infty}_{i=1}\frac{2W^{i}_{u_{1},u_{n_{5}+2}}(P_{n_{1}})}{\rho_{1}^{i}}\leq\frac{2}{\rho_{1}^{n_{5}+1}}(\frac{10\rho_{1}}{\rho_{1}-1}e^{\frac{2(n_{5}+1)}{\rho_{1}^{2}}}+\frac{32}{\rho_{1}-4}).$
Since $\rho_{1}\geq\max\left\{\sqrt{310},\sqrt{\sum_{1\leq i\leq 5}n_{i}}\right\}\geq\max\left\{\sqrt{310},\sqrt{5n_{5}+3}\right\}$, we have $e^{\frac{2(n_{5}+1)}{\rho_{1}^{2}}}=e^{\frac{(2n_{5}+\frac{6}{5})+\frac{4}{5}}{\rho_{1}^{2}}}<e^{0.403}<1.5$. Thus,
@@
$\leq\frac{2}{\rho_{1}^{n_{5}+1}}(\frac{10\rho_{1}}{\rho_{1}-1}e^{\frac{2(n_{5}+1)}{\rho_{1}^{2}}}+\frac{32}{\rho_{1}-4}-\rho_{1}-1)$
@@
Now consider $g(x)=14+\frac{15}{x-1}+\frac{32}{x-4}-x$ for $x\geq\sqrt{310}$. Clearly, $g(x)$ is decreasing for $x\geq\sqrt{310}$. By a calculation, $g(\sqrt{310})=-0.35\cdots$. Thus $g(x)<0$ for any $x\geq\sqrt{310}$. So, $\sum^{\infty}_{i=1}(\frac{W^{i}(G^{1})}{\rho_{1}^{i}}-\frac{W^{i}(G^{2})}{\rho_{1}^{i}})<0$, or equivalently, $\sum^{\infty}_{i=1}\frac{W^{i}(G^{1})}{\rho_{1}^{i}}<\sum^{\infty}_{i=1}\frac{W^{i}(G^{2})}{\rho_{1}^{i}}$. It follows that $f(G_{1},\rho_{1})>f(G_{2},\rho_{1})$. Note that $f(G_{1},\rho_{1})=1$. So, $f(G_{2},\rho_{1})<1$. Recall that $f(G_{2},x)$ is strictly increasing in its domain, and $f(G_{2},\rho_{2})=1$. We must have $\rho_{1}<\rho_{2}$. This completes the proof. $\Box$
@@
Let $G_{3}$ be the graph obtained from $G$ by adding the edge $u^{\prime}u^{\prime\prime}$ crossing the face $u^{\prime}u_{i_{0}}u^{\prime\prime}u_{i_{0}+1}u^{\prime}$. Clearly, $G_{3}$ is planar and $C_{\ell}$-free. But $\rho(G_{3})>\rho(G)$ by Lemma 2.2, contradicting that $G\in{\rm SPEX}(\mathcal{G}_{n,\ell})$. Therefore, $u^{\prime}u^{\prime\prime}\in E(G)$.
@@
Note. In this version, we correct some calculations, and include the detailed proof of Theorem 1.4.

### v3

@@ 2 Preliminaries
[…]
Now consider $g(x)=40+\frac{25}{x-1}+\frac{128}{x-8}-x$ for $x\geq\sqrt{2000}$. Clearly, $g(x)$ is decreasing for $x\geq\sqrt{2000}$. By a calculation, $g(\sqrt{2000})=-0.66\cdots<0$. Thus $g(x)<0$ for any $x\geq\sqrt{2000}$. Then $\sum^{\infty}_{i=1}(\frac{W^{i}(G^{1})}{\rho_{1}^{i}}-\frac{W^{i}(G^{2})}{\rho_{1}^{i}})<0$, or equivalently, $\sum^{\infty}_{i=1}\frac{W^{i}(G^{1})}{\rho_{1}^{i}}<\sum^{\infty}_{i=1}\frac{W^{i}(G^{2})}{\rho_{1}^{i}}$. It follows that $f(G_{1},\rho_{1})>f(G_{2},\rho_{1})$. Note that $f(G_{1},\rho_{1})=1$. So, $f(G_{2},\rho_{1})<1$. Recall that $f(G_{2},x)$ is strictly increasing in its domain, and $f(G_{2},\rho_{2})=1$. We must have $\rho_{1}<\rho_{2}$. This completes the proof. $\Box$
@@
$\rho_{j}\geq\rho(K_{1}\vee\overline{K_{|T|+\sum_{1\leq i\leq 5}n_{i}}})=\sqrt{|T|+\sum_{1\leq i\leq 5}n_{i}}\geq\max\left\{\sqrt{6500},\sqrt{\sum_{1\leq i\leq 5}n_{i}}\right\}.$
@@
$\sum^{\infty}_{i=1}\frac{2W^{i}_{u_{1},u_{n_{5}+2}}(P_{n_{1}})}{\rho_{1}^{i}}\leq\frac{2}{\rho_{1}^{n_{5}+1}}(\frac{30\rho_{1}}{\rho_{1}-1}e^{\frac{\frac{3}{2}(n_{5}+1)}{\rho_{1}^{2}}}+\frac{32\rho_{1}}{\rho_{1}-8}).$
Since $\rho_{1}\geq\max\left\{\sqrt{6500},\sqrt{\sum_{1\leq i\leq 5}n_{i}}\right\}\geq\max\left\{\sqrt{6500},\sqrt{5n_{5}+3}\right\}$, we have $e^{\frac{\frac{3}{2}(n_{5}+1)}{\rho_{1}^{2}}}\leq e^{\frac{\frac{3}{2}(n_{5}+1)}{5n_{5}+3}}<e^{0.4}<1.5$. Thus,
@@
$\leq\frac{2}{\rho_{1}^{n_{5}+1}}(\frac{30\rho_{1}}{\rho_{1}-1}e^{\frac{\frac{3}{2}(n_{5}+1)}{\rho_{1}^{2}}}+\frac{32\rho_{1}}{\rho_{1}-8}-\rho_{1}-1)$
@@
[…]
Now consider $g(x)=76+\frac{45}{x-1}+\frac{256}{x-8}-x$ for $x\geq\sqrt{6500}$. Clearly, $g(x)$ is decreasing for $x\geq\sqrt{6500}$. By a calculation, $g(\sqrt{6500})=-0.53\cdots$. Thus $g(x)<0$ for any $x\geq\sqrt{6500}$. So, $\sum^{\infty}_{i=1}(\frac{W^{i}(G^{1})}{\rho_{1}^{i}}-\frac{W^{i}(G^{2})}{\rho_{1}^{i}})<0$, or equivalently, $\sum^{\infty}_{i=1}\frac{W^{i}(G^{1})}{\rho_{1}^{i}}<\sum^{\infty}_{i=1}\frac{W^{i}(G^{2})}{\rho_{1}^{i}}$. It follows that $f(G_{1},\rho_{1})>f(G_{2},\rho_{1})$. Note that $f(G_{1},\rho_{1})=1$. So, $f(G_{2},\rho_{1})<1$. Recall that $f(G_{2},x)$ is strictly increasing in its domain, and $f(G_{2},\rho_{2})=1$. We must have $\rho_{1}<\rho_{2}$. This completes the proof. $\Box$
@@
Let $G_{3}$ be the graph obtained from $G$ by adding the edge $u^{\prime}u^{\prime\prime}$ crossing the face $u^{\prime}u_{i_{0}}u^{\prime\prime}u_{i_{0}+1}u^{\prime}$. Clearly, $G_{3}$ is planar and $C_{\ell}$-free. But $\rho(G_{3})>\rho(G)$ by Lemma 2.2, contradicting that $G\in{\rm SPEX}(\mathcal{G}_{n,\ell})$. Therefore, $u^{\prime}u^{\prime\prime}\in E(G)$. $\Box$